A car battery with a 12-V emf and an internal resistance of 0.050Ω is being charged with a current of 60 A. Note that in this process, the battery is being charged.

(a) What is the potential difference across its terminals? _____ V

(b) At what rate is thermal energy being dissipated in the battery? _____ W

(c) At what rate is electric energy being converted into chemical energy? _____ W


1
Expert's answer
2021-05-21T11:35:55-0400

a.) We have "e = 12V"


Internal resistance of the battery "= 0.05 \Omega"

Current "= 60A"

Terminal voltage of the battery "V = e-Ir"

The battery is being charged so, I is now negative:


"V = e + Ir"

"V = 12+(60 \times 0.05) = 15V"

b.) "P_{thermal} = I^2R = (60)^2 \times 0.05 = 180W"

c.) "P = \dfrac{E}{t} = IV"

"P = 60 \times 12 = 720W"

"P_{chemical} = 720-180 = 540W"